I have a question about linear systems using substitution.
How do solve:
x=6-4y
and
2x-3y=1

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

How do solve:
x=6-4y
and
2x-3y=1
------------
Sub for x in the 2nd eqn:
2(6-4y) - 3y = 1
Solve for y.
Then sub for y into either eqn and find x.

Answer by stanbon(75887)

systems using substitution.
How do solve:
x=6-4y
and
2x-3y=1
----
Substitute for "x" and solve for "y":
2(6-4y)-3y = 1
12-8y-3y = 1
12-11y = 1
-11y = -11
y = 1
---
Now solve for "x":
x = 6-4y
x = 6-4*1
x = 2
==========
Cheers,
Stan H.
===========